OpenStudy (anonymous):

(1-tanx)^2=sec^2 x-2 tanx

OpenStudy (anonymous):

ok!! now the left hand side \[(1-tanx)^2 =1-2tanx+tan^2x=(1+tan^2x)-2tanx=sec^2x-2tanx\]

OpenStudy (anonymous):

I go from the left hand side to the right hand side, that is the whole proof

OpenStudy (anonymous):

I still dont understand the second part

OpenStudy (anonymous):

which one?

OpenStudy (anonymous):

how do you get the 1-2tanx+tan^2x

OpenStudy (anonymous):

do you know \((a-b)^2 = a^2 -2ab +b^2\)

OpenStudy (anonymous):

your a is 1 your b is tan x just apply that expression

OpenStudy (anonymous):

i wasnt taught that hold on let me think about it

OpenStudy (anonymous):

Whaaaaat is wrong with your prof???

OpenStudy (anonymous):

a, b can be anything!!! for trig, for number, for letter for something else

OpenStudy (anonymous):

it's Virtual school so they miss things sometimes

OpenStudy (anonymous):

ok, take your time to think about it. It's 1:00 am now. I go to bed now. Good luck

OpenStudy (anonymous):

thank you so much! sorry for the bother... good night

OpenStudy (anonymous):

omg i got it you're AWESOME!!!!!!!!!!!!!!!!!!!!!!!!